Vocaloid (vocaloid):

Second Order Differential Equation with Non-constant Coefficients (question will be typed below)

Vocaloid (vocaloid):

\[\frac{ d^2y }{ dx^2 }-\frac{ dy }{ dx }-(\frac{ 6 }{ x^2 }+\frac{ 2 }{ x })y = 0\]

Vocaloid (vocaloid):

I think I'm supposed to use reduction of order but I'm not exactly sure where to start

OpenStudy (dumbcow):

Here are some notes/examples for reduction of order http://tutorial.math.lamar.edu/Classes/DE/ReductionofOrder.aspx

OpenStudy (dumbcow):

I believe you will need to start with an initial solution of the form: \[y = ax^2 + c\]

Vocaloid (vocaloid):

The problem says I'm looking for a value of lambda such that y = x^lambda is a solutino

Vocaloid (vocaloid):

If I'm not given an initial solution how do I find one?

OpenStudy (dumbcow):

yeah thats the tricky part, usually by an educated guess then checking to see if it works. since the non-constant coefficient has an x^2 in denominator, it makes sense for y to have an x^2 term. Also this allows for 1st derivative to be linear which will cancel out the "2y/x" part of coeffecient.

Vocaloid (vocaloid):

thank you, that's the part I was having some trouble with I'll give it a shot and see how it goes

OpenStudy (dumbcow):

good luck, do you want me to post the solution or you can check answer on wolfram too ?

Vocaloid (vocaloid):

I tried wolfram but it ended up giving me a pretty complex solution

OpenStudy (dumbcow):

looks like i was wrong then. Initial solution should be of the form y = c/x^2 or y= x^-2 Then try to get 2nd solution by solving for function v(x) y = v*x^-2

Vocaloid (vocaloid):

I got the answer, thank you for your help! I did a substitution of y = x^lambda in the original equation